The number of patients seen in the Emergency Room in any given hour is a random variable represented by x. The probability distribution for x is :

        X

      10

       11

      12

      13

      14

     P(X)

     0.4

      0.2

     0.2

     0.1

     0.1


1
Expert's answer
2022-02-24T06:32:12-0500

1.


"E(X)=\sum _ix_ip(x_i)=10(0.4)+11(0.2)"

"+12(0.2)+13(0.1)+14(0.1)=11.3"

"E(X^2)=\sum _ix_i^2p(x_i)=10^2(0.4)+11^2(0.2)"

"+12^2(0.2)+13^2(0.1)+14^2(0.1)=129.5"

"Var(X)=\sigma^2=E(X^2)-(E(X))^2"

"=129.5-(11.3)^2=1.81"

"\sigma=\sqrt{\sigma^2}=\sqrt{1.81}=1.3454"



2.

"P(X=14)=0.1""P(X\ge 14)=0.1"

"P(X\ge 12)=P(X=12)+P(X=13)"

"+P(X=14)=0.2+0.1+0.1=0.4"

"P(X\le 11)=P(X=10)+P(X=11)"

"=0.4+0.2=0.6"